OpenStudy (anonymous):

I need help with Algebra II The directions say "Simplify each expression. Assume that all variables are nonzero." (X^3)^-2

OpenStudy (anonymous):

yay so........................ how do you do this? ;) x^-y = 1/(x^y)

OpenStudy (anonymous):

@Luis_Rivera ... or we could ask him to do it -.-

OpenStudy (anonymous):

@Luis_Rivera if you give a man a fish he will live for a day, if you teach a man how to fish he will live for a lifetime, e.g. teach them, not tell them

OpenStudy (anonymous):

they're here to learn not to get help, iirc

OpenStudy (anonymous):

so you put one over the expression and multiply the exponents?

OpenStudy (anonymous):

@Luis_Rivera touche . also i misquoted slightly :(

OpenStudy (anonymous):

@tessa the point is 1/x = x^(-1)

OpenStudy (anonymous):

simiarly 1/(y^n) = y^(-n)

OpenStudy (anonymous):

Oh, so would the point be 1/x for any problem like this?

OpenStudy (amistre64):

think of a negative exponent as being underground.

OpenStudy (amistre64):

\[x^{-n}=\frac1{x^n}\]

OpenStudy (amistre64):

but thats just a cue; becuase they also give you things like:\[\frac{1}{x^{-n}}\]which we have to get it back up \[\frac{1}{x^{-n}}=x^n\]

OpenStudy (amistre64):

in short, negative exponents are on the wrong side of the tracks :)

OpenStudy (anonymous):

Thank you :)